# The Wisdom of Intellectually Humble Networks

### v1

@@ Introduction
Intuitively, in addition to estimation accuracy, people’s psycho-social attributes can also modulate how much weight they assign to their own estimates. In politically charged contexts, people’s extraordinary preference for cognitive consistency means they disproportionately accept information from ‘in-group’ members who already share similar beliefs (?, ?, ?, ?). Such partisan biases can persist even when people are offered financial incentives for accurate estimates (?, ?). In this work, we are interested in how the understudied yet highly promising psychological construct of intellectual humility modulates the weights people assign to their own (and peers’) estimates and how that process, in turn, affects the collective wisdom of social networks.
@@
where, $w_{1}^{(c)}$ and $w_{2}^{(c)}$ are the intercept and positive slope of the linear relation, respectively. $w_{2}^{(c)}$ is steeper for treatment networks than control networks to capture the superior evidence evaluation capabilities of the treatment agents.
@@
where, $w_{3}^{(c)}$ and $w_{4}^{(c)}$ are the intercept and negative slope of the linear relation, respectively. $w_{4}^{(c)}$ is less steep for treatment networks than control networks to capture the superior ability of the treatment agents to ignore partisan bias.
@@
We clip $\tilde{\alpha}_{i}^{(c)}$ to bind it between 00 and $1$, which has been observed to hold for nearly all users in our dataset.
@@ Discussion
This work contributes to the body of collective wisdom research, elucidating whether and how the psychological construct of intellectual humility can affect the errors and polarization of collectively held beliefs. We construct a simple model for encoding IH in the self-weight parameters of the DeGroot local information aggregation model. Using this model, we show that estimation errors and polarization are reduced in IH-enhanced networks and that the results are robust to different task contexts and network structures. Our simulations capture how IH fosters balanced belief revisions: less accurate individuals revise more effectively in response to social feedback, while more accurate individuals better retain their performance. This effect is showcased by an elevated Revision Coefficient in the IH-enhanced networks, suggesting a mechanism driving the collective performance benefits.
Our results corroborate and extend existing IH research that hypothesizes higher accuracy and reduced extremism at an individual level (?, ?, ?, ?, ?). We demonstrate that in closed social systems, the benefits of reduced errors and polarization should be readily observable if an empirical experiment can be designed with a practical IH intervention for the treatment condition.
However, an effective and scalable intervention for elevating IH in a network setting is not yet established. If such an intervention can be designed and empirical data can be collected from a treatment network, our simple model for encoding IH in the self-weight parameters can be validated and extended. For instance, our analytical model captures two core elements of IH (evidence quality evaluation and homophily-based bias resistance), but empirical data can allow us to capture further nuances of IH in a data-driven manner. Designing the scalable intervention and validating our model with empirical data comprise part of our future work. This research direction holds immense potential for practical application development in organizational decision-making, public policy, and collaborative problem-solving scenarios—not to mention in politically charged contexts for combating collective inaccuracies in human societies.

### v2

@@ Introduction
Intuitively, in addition to estimation accuracy, people’s psycho-social attributes can also modulate how much weight they assign to their own (and peers’) estimates. In politically charged contexts, people’s extraordinary preference for cognitive consistency means they disproportionately accept information from ‘in-group’ members who already share similar beliefs (?, ?, ?). Such partisan biases can persist even when people are offered financial incentives for accurate estimates (?, ?). In this work, we are interested in how the understudied yet highly promising psychological construct of intellectual humility modulates the weights people assign to their own (and peers’) estimates and how that process, in turn, affects the collective wisdom of social networks.
@@
where, $w_{1}^{(c)}$ is the intercept and $w_{2}^{(c)}$ is the positive slope of the linear relation. $w_{2}^{(c)}$ is steeper for treatment networks than control networks to capture the superior evidence evaluation capabilities of the treatment agents.
@@
where, $w_{3}^{(c)}$ is the intercept and $w_{4}^{(c)}$ is the negative slope of the linear relation. $w_{4}^{(c)}$ is less steep for treatment networks than control networks to capture the superior ability of the treatment agents to ignore partisan bias.
@@
We clip $\tilde{\alpha}_{i}^{(c)}$ to bind it between 00 and $1$, a property observed to hold for nearly all users in our dataset.
@@ Discussion
This work contributes to the body of collective wisdom research, illustrating whether and how the psychological construct of intellectual humility can affect the errors and polarization of collectively held beliefs. Our model encodes IH in the self-weight parameters of the DeGroot process. Using this model, we show that IH-enhanced networks exhibit lower errors and polarization, with robust results across tasks and network structures. Our simulations capture how IH fosters balanced belief revisions: less accurate individuals revise more effectively, while more accurate individuals better retain their performance. This effect is showcased by an elevated Revision Coefficient in the IH-enhanced networks, suggesting a mechanism driving the collective performance benefits.
A body of agent-based simulations previously probed how psycho-social elements (e.g., mutual trust) and network structural elements (e.g., network size and connectivity) can modulate polarization, echo chambers, and inaccuracies in epistemic interaction-based social networks (?, ?, ?, ?, ?). For instance, when Bayesian agents become trusting of their peers’ credibility, a snowball effect can ensue that polarizes even rational agents (?, ?, ?). Moreover, network structural properties can aid echo chamber formation, while cognitive and social biases can exacerbate the process (?, ?, ?). Our work augments this research thread by showing that IH can reduce polarization and collective errors by encouraging balanced belief revisions, and that the effects remain robust across network structures.
Our results also corroborate and extend existing IH research that hypothesizes higher accuracy and reduced extremism at the individual level (?, ?, ?, ?). We show that these benefits can manifest in closed social systems if the interactants’ IH can be practically elevated. However, designing a scalable intervention for elevating IH in a network setting is non-trivial and formulates part of our future work. Such empirical data can help validate and extend our model by elucidating further nuances of IH. This research direction holds immense potential for informing applications in organizational decision-making, public policy, and collaborative problem-solving—not to mention in politically charged contexts for combating collective inaccuracies in societies.
Code Availability